OpenStudy (mackenzie2013):

The cost to mail a letter is a base charge of $0.42 plus $0.17 for each ounce. Write an equation that could be used to find how many ounces a letter that costs $0.93 weighs. (Let x represent the number of ounces.)

OpenStudy (mackenzie2013):

0.59x = 0.93 0.42x + 0.17 = 0.93 0.59 = 0.93x 0.42 + 0.17x = 0.93

OpenStudy (whpalmer4):

the cost is the base charge ($0.42) plus the per-ounce cost ($0.17) * number of ounces (x) cost = 0.42 + 0.17*x we know the cost is $0.93 set those two equal to each other